OpenStudy (anonymous):

Why is the derivative of (1-x)^3 negative? I can see 3(1-x)^2, but why -3(1-x)^2?

OpenStudy (anonymous):

That's what the chain rule tells us. In other words, that is because the derivative of (1-x) is -1.

OpenStudy (anonymous):

Thanks, AnwarA.

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

(U^n ) = n U^(n-1) * U' U = 1 -x -> U' = -1 => derivative result is negative!